WMT Corporation has an issue of common stock. The firm most recently paid a dividend of $1.00 per share. The dividend is projected to grow at the rate of 25% for the next two years followed by 6% thereafter for indefinite future. If the investor demand 12% return on similar risk stocks, what should be the price of this common stock?
Answer #1

D1=(1*1.25)=1.25

D2=(1.25*1.25)=1.5625

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(1.5625*1.06)/(0.12-0.06)

=27.6041667

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.25/1.12+1.5625/1.12^2+27.6041667/1.12^2

=$24.37(Approx).